OpenStudy (anonymous):

Use the definition of continuity to determine whether f is continuous at a. f(x) = x - 4/x +5  a = 4 A. Not continuous B. Continuous

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

f(x) will continous at point x = a if the value of f(a) = limit (x->a) f(x) let's check it : f(x) = (x - 4)/(x +5), so the value of f(4) = f(4) = (4-4)/(4+5) = 0/9 = 0 lim (x->4) f(x) = lim (x->4) (x - 4)/(x +5) = (4-4)/(4+5) = 0/9 = 0 because they are same, therefore f(x) be continous at point x = 4

OpenStudy (anonymous):

wow thank again @RadEn

OpenStudy (raden):

you're welcome :)

OpenStudy (anonymous):

;-)